OpenStudy (anonymous):

Let x-2y-z=-2 and -2x+4y+z=5 be two planes. Find the parametric equations for the line in which x-2y-z=-2 and -2x+4y+z=5 intersect.

OpenStudy (anonymous):

Anybody can help me?!

OpenStudy (zarkon):

find the normal vectors for the two planes find the cross product then find any point that lies in both planes form equation I get \[(2,1,0)t+(-3,0,-1)\]

OpenStudy (anonymous):

how to find normal vectors?

OpenStudy (zarkon):

take the coefficients from the equations of the planes x-2y-z=-2 and -2x+4y+z=5 (1,-2,-1) and (-2,4,1)